OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. A 40 foot beam is to be cut in two. If one portion is one third the length of the other portion, what are the two lengths of the portions? please show work!

OpenStudy (anonymous):

1) one part=y; another=y 2) x+y=40 3) x=y/3 so, per 1) you'll have: 4) y+y/3 = 40 solve for y: 4y/3 = 40 5) y=30 now, substitute value of y in 3) you'll have x=10 6) check: x+y=10+30=40 True!